OpenStudy (anonymous):

Problem: Two particles are moving in the xy-plane. The move along straight lines at constant speed. At time t, particle A's position is given by x=t+2, y=1/2t-3 and particle B's position is given by x=12-2t, y=6-1/3t. (a) Find the equation of the line along which particle A moves. (b) Find the equation of the line along which particle B moves. (c) Find the time (value of t) at which the distance between A and B is minimal. NEED HELP PLEASE.. :)

OpenStudy (anonymous):

Okay, so when they say "Fine the equation of the line" I think it to mean they want you to find \(y = f(x)\). If that is the case then my solution would be: 1) Solve for t in terms of x. \(t = f(x)\) 2) Plug it into y. \(y = g(t), f(x) =g(t(x))\).

OpenStudy (anonymous):

those parametric equations don't describe straight lines...

OpenStudy (anonymous):

For (c) they seem to want you to find the minimum euclidean distance. So I would just plug the equations into the distance formula and use the calculus you learned for minimization (find critical numbers, etc).\[ d(t) = \sqrt{[x_1(t) - x_2(t)]^2+[y_1(t) - y_2(t)]^2} \]

OpenStudy (anonymous):

@dpaInc Yeah I was wondering about that too.... maybe they meant curves?

OpenStudy (anonymous):

possibly...

OpenStudy (anonymous):

the equation of the line along which A moves y=(1/2)x-5 the equation of the line along which B moves y=(1/6)x+12 the time at which the distance is minimal happens when both particles are at (-36,6)